the following equations describe the cost of production and demand for soccer balls.
demand price = 10 - q
MR = 10-2q
TC = 3+q+0.5q^2
MC = 1+Q
I think i need to figure out what Q is to solve these.. MR i think i'd set the equation and i'd end up with =5 I need to find what the fixed cost is for the soccer balls and how many are produced to in the econoamy if they profit max. what price do each sell for and what is the profit earned.. any help is appreiciated thanks
1
Expert's answer
2012-05-17T07:44:29-0400
if they profit max, MR = MC = P, 10 - 2Q = 1 + Q, Q = 3, P = 4, FC is a free number from the equation of
TC and is equal 3. TC = 3 + 3 + 0.5*3^2 = 10.5. Profit will be TP = TR - TC =
P*Q - TC = 4*3 - 10.5 = 1.5